# Comparative study of quantum emitter fabrication in wide bandgap materials using localized electron irradiation

### 07 Dec 2023 v2

@@ Abstract
Quantum light sources are crucial foundational components for various quantum technology applications. With the rapid development of quantum technology, there has been a growing demand for materials that are capable of hosting quantum emitters. One such material platform are fluorescent defects in hexagonal boron nitride (hBN) inducing deep sub-levels within the band gap. The question arises if other layered wide bandgap (2D) materials offer similar single photon emitting defects. Here, we investigate and compare the fabrication of quantum emitters in exfoliated multi-layer mica flakes with hBN and other wide bandgap 3D crystals (silicon carbide and gallium nitride) which are known to host quantum emitters. We use our primary fabrication technique of localized electron irradiation using a standard scanning electron microscope. To complement our experimental work, we employ density functional theory simulations to study the atomic structures of intrinsic defects and their photophysical properties. While our fabrication technique can create hBN quantum emitters with a high yield and high single photon purity, it is unable to fabricate emitters in the other solid-state crystals under investigation. This allows us to draw conclusions on the emitter fabrication mechanism, which could be relying on the activation of already present defects by charge state manipulation. We therefore provide an important step toward the identification of hBN emitters and their formation process.
@@ I Introduction
The quantum emitters can be fabricated using various methods such as strain activation using nanostructure [24], mechanical damage using the tip of an atomic force microscope (AFM) [28], thermal activation of naturally occurring defects [7], plasma [29] and chemical etching [30], as well as energetic irradiation [25, 31, 32, 33, 34]. Of particular interest is the formation of emitters at pre-defined locations, while minimizing the impact of the crystal environment, for which localized irradiation has shown to be useful. The response of materials to the irradiation by energetic particles such as electrons and ions has been studied intensively [35]. Irradiation can create new defects in solids by knocking out atoms or implanting impurities, but also modify or activate existing defects by breaking and re-organizing bonds or changing the charge state. This can ultimately result in the modification of their physical and chemical properties. Recent works have shown the high yield of quantum emitters in TMDs and hBN by modifying the surface properties or doping [33, 32, 25, 34] and arrays of photophysically identical emitters with correlated dipoles [36], all using electron irradiation of hBN or a focused helium beam in case of MoS.
The emitter formation process using electron irradiation in hBN is not well understood. Not only differ the emitters in terms of formation depth [26] compared to other fabrication methods, but also the emitter properties differ when comparing the photophysical properties of electron beam irradiated hBN emitters in different works: some have basically random-oriented dipoles [37] while others have correlated dipole directions [36]; or some emit blue light [33, 32], and others have their optical transition in the yellow region [34, 36]. The question arises how unique the electron irradiation process is to hBN or if it can be applied to other wide bandgap materials as well.
@@
In this work, we carry out a comprehensive study and compare our electron irradiation fabrication method on hBN [32, 33, 34] with other wide bandgap materials, including 2D mica muscovite with a bandgap around 5 eV [38, 39], as well as the 3D crystals silicon carbide and gallium nitride. Both the latter are well known to host optically active quantum emitters [9, 10, 40, 8, 11, 41], while in the case of mica, only fluorescent molecules adhering to the surface have been shown to support quantum emission [42]. By using different materials and crystals (of the same type) but from different vendors we gain insight into the emitter formation process. We first fabricate an array of high quality quantum emitters in hBN and then replicate the same fabrication process on a large variety of samples. In addition, we also try high temperature annealing and check if any newly created defects require thermal activation. Our experiments and conclusions are supported using density functional theory (DFT) calculations, which reveal the electronic structure of potential defects in mica and allows us to infer if the defect should be optically active or not. Finally, we draw conclusions about the quantum emitter formation process in hBN.
II Experimental details
II.1 Sample preparation
@@
To create hBN emitters, we first mechanically exfoliate from bulk crystal onto an organic polymer sheet made of PDMS (polydimethylsiloxane) using a scotch tape. Afterwards, a suitable thin flake is transfer onto a grid patterned standard silicon substrate with a 290 nm thermal oxide layer by dry stamping (see Methods). We have used bulk hBN from HQ Graphene (named hBN-1 in the following) and 2D Semiconductors (hBN-2), which produce their crystals using different processes, e.g., 2D Semiconductors utilize a high-pressure anvil cell growth method. It is therefore likely that the samples have a different intrinsic doping of impurities.
@@
The mica muscovite flakes are exfoliated and transferred in the same fashion as mentioned above. Fig. 1(a) shows an optical image of a transferred mica flake. The thickness map is created using a phase-shift interferometer (PSI) that measures the optical path length (OPL) through the flake and is shown in Fig. 1(b). To convert the OPL measurements into physical thickness values, rigorous coupled-wave analysis (RCWA) simulations are used, similarly as it has been done before for TMDs [43] and hBN [29]. The RCWA model shown in Fig. 1(c) is calibrated using real thickness measurements obtained with an AFM (see Supplementary Section S1). Once the RCWA model is calibrated, using the PSI proves to be both accurate and faster than the commonly employed AFM thickness measurements [43, 29, 26]. It is important to note, however, that this technique can only yield accurate results for OPL thicknesses up to 30 nm (for mica on silicon substrate with a 290 nm thermal oxide layer), as the relation between physical and optical thickness follows an S-curve due to interference (see Fig. 1(c)). The physical flake thickness through the dashed line in the OPL map is shown in Fig. 1(d). The flake on which the subsequent electron irradiation is carried out has a thickness around 6 nm (blue area in the microscope image) and around 10 nm (teal area in the microscope image).
@@ II.2 Electron irradiation
We performed localized electron irradiation using a SEM to deliberately introduce defect-based quantum emitters into the materials. This method has demonstrated its effectiveness in creating quantum emitters within hBN, as substantiated by prior studies [32, 33, 34]. The high lateral resolution of SEM allow us to irradiate the exfoliated multi-layer mica flake at a chosen spot as shown in Figure 1(e). The irradiated spots are clearly visible in the resulting SEM image post irradiation process. Here each spot is irradiated using an accelerating voltage of 3 kV and electron current of 25 pA with a dwell time of 10 s (see Methods). This resulted an electron fluence of cm at each irradiated spot. Note that we used low electron fluence (10 cm) is used to record the image of the flake in order to avoid the creation of randomly located emitters in the flake, due to electron beam scanning during imaging process.
@@
To control the electron dose precisely, we adjust the irradiation time during the irradiation process. Additionally, we performed the irradiation process with varying electron irradiation times on one of the flakes. The resulting PL map of the irradiated spot within the flake is displayed in the inset of Figure 1(f). To calculate the average PL intensity, we took the mean across the same number of pixels for each spot (see Supplementary Section S2). The observed behavior is quite intriguing; the average PL intensity exhibits saturation tendencies with increasing irradiation time, as depicted in Figure 1(f). This behavior suggests that the creation of defects in the mica flake is constrained by an upper limit, resembling the controlled creation of emitters in hBN through the irradiation process [44, 34]. These findings underscore the precision and controllability of defect engineering within the mica flakes, promising exciting prospects for applications in quantum emitters fabrication.
III Results and discussions
III.1 Optical characterization
@@
To investigate our fabrication methodology of quantum emitters, we optically characterized the samples using a commercial fluorescence lifetime imaging microscope (see Methods). We first investigated irradiated mica flakes. We performed all the measurements at room temperature using different pulsed excitation laser with a repetition rate of 20 MHz. We always excite below bandgap to exclusively excite any states that were created during the irradiation process and at the same time to avoid the bandgap excitation of the free exciton. Figures 2(a)-(d) present the PL maps of the samples using 375 nm, 470 nm, 530 nm and 640 nm pulsed excitation lasers, respectively. Under 375 nm excitation laser, we observe weak emission from the irradiated spots indicating very inefficient excitation in the UV region. However, we notice bright emission from the irradiated spots under different excitation lasers as shown in Figure 2(b)-(d). Bright emission in the PL map corresponds to the irradiated spots as can be confirmed by SEM and PL images. We also did not observe any correlation in between bright emission from irradiated spots and the flake thickness.
@@
Furthermore, we see diffraction-limited isolated spots in the close vicinity of the irradiated spots as shown in Figure 1(e), when a zoomed-in PL maps is created using 470 nm pulsed excitation laser. This is similar to the previously studied hBN irradiation experiments, which were responsible for the generation of blue and yellow single photon emitters around the irradiated spots [32, 33, 34, 36]. However, we did not find any sharp spectral feature as recorded under 470 nm excitation laser in combination with a long pass filter with a cut-off wavelength at 500 nm to suppress the excitation laser. We rather observe a broad peak centered around 600 nm as shown in Figure 2(f). This emission behaviour resembles what we observed from the irradiated spots in hBN in our previous works [34, 36]. This could be an indication that our irradiation process may indeed lead to the formation of similar defect complexes in mica as well upon electron irradiation, potentially due to deposited carbon [34]. Note that due to the relatively weak PL emission, we extended the exposure time to acquire the spectrum effectively. As a result of this longer exposure time, we also captured certain experimental artifact peaks in our measurements. Notably, these artifact peaks were also present in our background spectrum measurements, which were taken at a location away from the irradiated spot and subjected to the same extended exposure time (see Supplementary Section S3).
@@
In order to further investigate the quantum nature of emission, we measure the second-order correlation function (g) from the various irradiated spots. In Figure 2(e), we present experimentally measured g curve, acquired using a 470 nm pulsed excitation laser with a repetition rate of 20 MHz. To analyze the data, we employed a model function to fit the g curve, and extracted a g around 0.86(5), which is larger than 0.5 (the generally accepted criterion for single photon emission due to a non-zero overlap with the single photon Fock state). Unlike a previous recent studies which introduced organic fluorescent molecules from solvents [42], we have not been able to create single emitters in mica using our electron irradiation process. We have also measured the excited state lifetime of the emitter to be around 1.2 ns. The lifetime decay curve is fitted with bi-exponential function to extract the lifetime as shown in Figure 2(h). Note that the initial sharp peak is due to instrument response function and taken care of in our fitting routing. Additionally, We examined the intensity from the irradiated spot as a function of excitation laser power, revealing the typical saturation curve for two level system, as depicted in Figure 2(i). We fitted the experimental data with model function revealing the average saturation power to be around 40.75 W and saturation intensity is 35.18 counts per second. Remarkably, our experimental observation did not reveal any photo-bleaching or blinking. effects under pulsed excitation. We measured the photon count rate over an extended time period as shown in Figure 2(j), and this revealed a stable emission with a stability of 11.4%.
@@
To gain deeper insights into the generality of our fabrication methodology in generating quantum emitters, we extended our irradiation experiment on various other insulating materials including two-dimensional hBN-1, hBN-2, silicon carbide (4H-SiC) and gallium nitride (GaN) crystals. For hBN-1, we initially exfoliated a multi-layer flake onto a standard Si/SiO substrate and conducted localized irradiation at a chosen spot within the flake, employing a dwell time of 10 seconds. This resulted in a similar electron fluence as in the case of mica irradiation. The resulting SEM of the flake is shown in Figure 3(a) (see Methods). Subsequently, we recorded the PL map of the irradiated flake using 530 nm pulsed excitation laser as shown in Figure 3(b). It is evident that the bright spot in the PL map corresponds to the irradiated spots in the SEM images. Our irradiation process effectively yielded emitters emitting at 575 nm [34].
To further understand quantum nature of emission, we performed g measurements using 530 nm pulsed excitation laser with repetition rate of 20 MHz. The results, displayed in Figure 3(c), reveal a clear g dip with a g(0) value of 0.084(2), as extracted from the model fitting function. We also noticed a stable emission from the quantum emitters with no instances of blinking and photo-bleaching as observe in Figure 3(d). The further photo-physical properties of as fabricated emitters are also summarized in the our previous works [34, 36].
For hBN-2, we performed the localized irradiation with the same electron beam parameters using the same process on the same SEM. Figure 3(a) shows the SEM images of an irradiated sample. The irradiated spots are clearly visible as distinctive dots. The corresponding PL map, depicted in Figure 3(b), also reveals bright emissions emanating from the irradiated spots. The PL map is created using 530 nm pulsed excitation laser with the similar laser power. This indicate the generation of similar defect complex formation due to localized irradiation. The zoomed-in PL map around one of the irradiated spot also reveals the formation of isolated diffraction limited spot, indicating the formation of isolated emitter as observe for hBN-1. However, in the consecutive PL map in Figure 3(c) shows that emitter bleached out eventually.
Furthermore, we investigated other emitters around the irradiated spots and recorded the PL count rate over time. and found that these emitters are not stable once excited. We observed the PL counts drops down to the noise level of the detector within the 10 s (see Supplementary Section S3). Such instability has not been observed for the emitters in hBN-1 under pulsed excitation even after prolonged exposure. Notably, we also observed no aging effect on the emitters in hBN-1. This instability observed in hBN-2 could be attributed to unstable charge carriers in the excited state potentially due to intrinsic doping of the material which depend on the growth conditions.
@@
In addition to the hBN, our irradiation experiment of SiC (4H-semi insulating) and GaN do not activate any emitters or emitters ensembles as evident by PL map (see Supplementary Section S4). We found that this is independent of the electron beam dose and electron accelerating voltage. We can draw already two summarizing conclusions from our experiments: (i) the electron irradiation process works very well to generate single photon emitters in hBN but not in other wide bandgap materials. This indicates that, at least in this fabrication method, lattice defects and not deposited complexes on the surface. If the latter was the case, this should also work on other materials as a substrate. And (ii), while we could create single photon emitters in both hBN samples, there seems to be a different intrinsic defect density that instabilizes the defects in hBN-2.
III.2 Theoretical calculations

### 26 Feb 2024 v4

@@ Abstract
Quantum light sources are crucial foundational components for various quantum technology applications. With the rapid development of quantum technology, there has been a growing demand for materials with the capability of hosting quantum emitters. One such material platform uses fluorescent defects in hexagonal boron nitride (hBN) that can host deep sub-levels within the band gap. The localized electron irradiation has shown its effectiveness in generating deep sub-levels to induce single emitters in hBN. The question is whether localized (electron beam) irradiation is a reliable tool for creating emitters in other wide bandgap materials and its uniqueness to hBN. Here, we investigate and compare the fabrication of quantum emitters in hBN and exfoliated muscovite mica flakes along with other 3D crystals, such as silicon carbide and gallium nitride, which are known to host quantum emitters. We use our primary fabrication technique of localized electron irradiation using a standard scanning electron microscope. To complement our experimental work, we employ density functional theory simulations to study the atomic structures of defects in mica. While our fabrication technique allows to create hBN quantum emitters with a high yield and high single photon purity, it is unable to fabricate single emitters in the other solid-state crystals under investigation. This allows us to draw conclusions on the emitter fabrication mechanism in hBN, which could rely on activating pre-existing defects by charge state manipulation. Therefore, we provide an essential step towards the identification of hBN emitters and their formation process.
@@ II Introduction
The quantum emitters can be fabricated using various methods such as strain activation using nanostructure [24], mechanical damage using the tip of an atomic force microscope (AFM) [28], thermal activation of naturally occurring defects [7], plasma [29] and chemical etching [30], as well as energetic irradiation [25, 31, 32, 33, 34, 35, 36]. Of particular interest is the formation of emitters at pre-defined locations, while minimizing the impact of the crystal environment, for which localized irradiation has shown to be useful. The response of materials to the irradiation by energetic particles such as electrons and ions has been studied intensively [37]. Irradiation can create new defects in solids by knocking out atoms or implanting impurities, but also modify or activate existing defects by breaking and re-organizing bonds or changing the charge state. This can ultimately result in the modification of their physical and chemical properties. Recent works have shown the high yield of quantum emitters in TMDs and hBN by modifying the surface properties or doping [35, 34, 25, 36] and arrays of photophysically identical emitters with correlated dipoles [38], all using electron irradiation of hBN or a focused helium beam in case of MoS [25].
The emitter formation process through electron irradiation in hBN remains inadequately understood. The emitters not only differ in terms of formation depth when compared to alternative methods such as plasma-activated emitters followed by thermal annealing [26] but also exhibit varying photo-physical properties. Some present randomly oriented dipoles [39], while other shows correlated dipole direction [38]. Additionally, there are some emitting blue light [35, 34], and others having their optical transitions in the yellow regions [36, 38].
While our previous work [36] focused on localized irradiation, which resulted in yellow emitters with high yield, it is still unable to address the uniqueness of the electron irradiation process to hBN and its applicability to other wide band gap materials. This is because, upon irradiation, we notice the formation of the surface complex (black dots in SEM images), which could be responsible for generating the emitters. Based on this idea, here, we explore how and whether our fabrication method can promise to generate emitters independent of material choice. For this, we consider two different types of hBN crystals grown differently, and analyze whether this strategy is applicable to different wide band gap materials. Moreover, with this scenario, we make use of the significant conclusion discussed in [35], in which it has been demonstrated that the generation of blue emitters shows a strong correlation with a sharp UV emission at 305 nm attributed to the presence of carbon dimers.
@@
More precisely, along with hBN, we considered mechanically exfoliated 2D muscovite mica with a band gap above 5 eV [40, 41] and 3D crystals such as silicon carbide and gallium nitride. While mica is not studied very well from the prospective to host quantum emitters, both the other 3D crystals are well known to host optically active quantum emitters [9, 10, 42, 8, 11, 43]. Furthermore, to understand the material growth sensitivity of electron irradiation, we utilize two different types of hBN crystals grown through different methods obtained from separate vendors. We first created an array of high-quality emitters in our standard hBN crystal and then replicated the same fabrication process on the other samples. In addition, we also try high-temperature annealing and check if any newly created defects require thermal activation. Our experiments and conclusions are supported using density functional theory (DFT) calculations, which reveal the electronic structure of potential defects in mica and allow us to infer whether the defect should be optically active.
III Experimental details
III.1 Sample preparation
@@
To create hBN emitters, we first mechanically exfoliate from bulk crystal onto an organic polymer sheet of PDMS (polydimethylsiloxane) using scotch tape. Afterwards, a suitable thin flake is transferred onto a grid-patterned standard Si/SiO substrate with a 290 nm thermal oxide layer by dry stamping (see Methods). We have used bulk hBN from HQ Graphene (named hBN-1 in the following) and 2D Semiconductors (hBN-2), which produce their crystals using different processes, e.g., 2D Semiconductors utilize a high-pressure anvil cell growth method. It is, therefore, likely that the samples have a different intrinsic doping of impurities.
@@
The muscovite mica flakes are exfoliated and transferred in the same fashion as mentioned above. Figure 1(a) shows an optical image of a transferred mica flake. The thickness map is created using a phase-shift interferometer (PSI) that measures the optical path length (OPL) through the flake and is shown in Figure 1(b). To determine the physical thickness of the OPL, we use Rigorous Coupled-Wave Analysis (RCWA) simulations. This method is similar to what has been previously done for TMDs [44] and hBN [29]. The RCWA model depicted in Figure 1(c) is calibrated using real thickness measurements obtained with an AFM (see Supplementary Section S1). Once the RCWA model is calibrated, the use of PSI proves to be both accurate and faster than the commonly employed AFM thickness measurements [44, 29, 26]. It is important to note, however, that this technique can only yield accurate results for OPL thicknesses up to 30 nm (for mica on Si/SiO substrate with a 290 nm thermal oxide layer), as the relation between physical and optical thickness follows an S-curve due to interference (see Figure 1(c)). The physical flake thickness through the dashed line in the OPL map is shown in Figure 1(d). The flake on which the subsequent electron irradiation is carried out has a thickness of around 6 nm (blue area in the microscope image) and around 10 nm (teal area in the microscope image).
@@ III.2 Electron irradiation
We performed localized electron irradiation using a scanning electron microscope (SEM) to fabricate defect-based quantum emitters into the materials. This method has demonstrated its effectiveness in creating quantum emitters within hBN, as substantiated by prior studies [34, 35, 36]. The high lateral resolution of SEM allows us to irradiate the sample within the limit of electron beam diameter. The irradiation is performed at a user selected spot within a pre-acquired SEM image (see Supplementary Section S2). Figure 1(e) presents an SEM image of a multi-layer mica flake post-irradiation, clearly showcasing the irradiation spots. Here, each spot is irradiated using an accelerating voltage of 3 kV and an electron current of 25 pA with a dwell time of 10 s (see Methods). While the theoretical resolution limit of SEM in field-free mode is around 10 nm, the used electron beam settings and beam alignment result in an electron beam diameter of around 300 nm. Since this is already the diffraction limit of our optical setup in subsequent measurements, we did not perform any further beam alignments. This resulted in an electron fluence of cm at each irradiated spot. Note that we used low electron fluence (10 cm) to record the image of the flake to avoid the creation of randomly located emitters due to electron beam scanning during the imaging process.
@@
To control the electron dose precisely, we adjusted the irradiation time during the irradiation process. Additionally, we performed the irradiation process with varying electron irradiation times on one of the flakes. The resulting PL map of the irradiated spot within the flake is displayed in the inset of Figure 1(f). The average PL intensity saturates with the irradiation time as shown in Figure 1(f), where average PL intensity is determined by averaging over the same number of pixels for each spot (see Supplementary Section S3). This behaviour suggests that defects in the mica flake are constrained by an upper limit, resembling the controlled creation of emitters in hBN through the irradiation process [45, 35, 36]. These findings underscore the precision and controllability of defect engineering within the mica flakes, promising exciting prospects for applications in quantum emitters fabrication.
IV Results and discussions
IV.1 Optical characterization
@@
To investigate our fabrication methodology of quantum emitters, we optically characterized the samples using a commercial fluorescence lifetime imaging microscope (see Methods). We first investigated irradiated mica flakes. We performed all the measurements at room temperature using different pulsed excitation lasers with a repetition rate of 20 MHz. We always excite below bandgap to exclusively excite any states that were created during the irradiation process and at the same time to avoid the excitation of the free exciton across the bandgap. Figures 2(a)-(d) present the PL maps of the irradiated mica flakes using 375 nm, 470 nm, 530 nm and 640 nm pulsed excitation lasers, respectively. Under 375 nm excitation laser, we observe weak emission from the irradiated spots indicating very inefficient excitation in the UV region. However, we notice bright emission from the irradiated spots under different excitation lasers as shown in Figure 2(b)-(d). Bright emission in the PL map corresponds to the irradiated spots as can be confirmed by SEM and PL images. We also did not observe any correlation in between bright emission from irradiated spots and the flake thickness.
@@
Furthermore, we see diffraction-limited isolated bright spots in the close vicinity of the irradiated spots, as shown in Figure 1(e), when a zoomed-in PL map is created using a 470 nm pulsed excitation laser. This is similar to the previously studied hBN irradiation experiments, which were responsible for the generation of blue and yellow single photon emitters around the irradiated spots [34, 35, 36, 38]. However, we did not find any sharp spectral feature resembling ZPL, as recorded using a 470 nm excitation laser in combination with a long-pass filter with a cut-off wavelength at 500 nm to suppress the excitation laser. We instead observe a broad peak centred around 600 nm as shown in Figure 2(f). The emission spectrum is similar to that observed from the irradiated spots in hBN [36, 38]. This observation could indicate that our irradiation process may indeed lead to the formation of similar defect complexes in mica, possibly induced by electron irradiation and potentially associated with deposited carbon. Note that we extended the exposure time to acquire the spectrum effectively due to the relatively weak PL emission. As a result of this longer exposure time, we also captured certain experimental artefact peaks in our measurements. Notably, these artefact peaks were also present in our background spectrum measurements, which were taken at a location away from the irradiated spot and subjected to the same extended exposure time (see Supplementary Section S4).
@@
To further investigate the quantum nature of emission, we measure the second-order correlation function (g) from the various irradiated spots. In Figure 2(e), we present an experimentally measured g curve, acquired using a 470 nm pulsed excitation laser with a repetition rate of 20 MHz. To analyze the data, we employed a model function to fit the g curve and extracted a g around 0.86(5), which is larger than 0.5 (the generally accepted criterion for single photon emission due to a non-zero overlap with the single photon Fock state). Unlike previous recent studies which introduced organic fluorescent molecules from solvents [46], we have not been able to create single emitters in mica using our electron irradiation process. The excited state lifetime of the emitters is estimated to be around 1.2 ns. The lifetime decay curve is fitted using a bi-exponential model function, as shown in Figure 2(h). It is worth noting that our fitting routine accounts for the instrument response function (IRF) shown in Figure 2(h), which has a relatively faster response time below 100 ps. Additionally, we examined the intensity from the irradiated spot as a function of excitation laser power, revealing the typical saturation curve for the two-level system, as depicted in Figure 2(i). We fitted the experimental data using a model function revealing the average saturation power to be around 40.75 W, and saturation intensity is 35.18 counts per second. The observed saturation count rate is significantly lower compared to the typical hBN emitters, including those induced by irradiation [34, 35, 36]. This implies that mica irradiated emitters have poor quantum efficiency, suggesting that mica may not be a suitable material for generating emitters through the irradiation process. Remarkably, our experimental observation did not reveal any photo-bleaching or blinking effects under pulsed excitation. We measured the photon count rate over an extended period as shown in Figure 2(j), which revealed a stable emission with a stability of 11.4%.
@@
To gain deeper insights into the generality of our fabrication methodology in generating quantum emitters, we extended our irradiation experiment on various other insulating materials, including hBN-1, hBN-2, silicon carbide (4H-SiC) and gallium nitride (GaN) crystals. For hBN-1, we initially exfoliated a multi-layer flake onto a standard Si/SiO substrate and conducted localized irradiation at a chosen spot within the flake, employing a dwell time of 10 seconds. This resulted in a similar electron fluence as in the case of mica irradiation (see Methods). Figure 3(a) shows the resulting SEM image of the flake. Subsequently, we recorded the PL map of the irradiated flake using a 530 nm pulsed excitation laser as shown in Figure 3(b). The bright spot in the PL map corresponds to the irradiated spots, as evident in the SEM image. This is consistent with our previous study that yields single emitters at 575 nm [36]. To further understand the quantum nature of emission, we performed g measurements using a 530 nm pulsed excitation laser with a repetition rate of 20 MHz. Figure 3(c), reveal a clear g dip indicating quantum nature of emission. The g(0) value of 0.084(2) is extracted from the model fitting function. We also noticed a stable emission from the quantum emitters with no instances of blinking and photo-bleaching, as observed in Figure 3(d). The further photo-physical properties of as-fabricated emitters are also summarized in our previous works [36, 38].
For hBN-2, we conducted irradiation using the same electron beam parameters (see Methods). The resulting SEM image of the flake is shown in the 3(e), and the corresponding PL map of the flake is presented in Figure 3(f), revealing bright emission from the irradiated spots. The PL map is generated using a 530 nm pulsed excitation laser with a similar average laser power (see Methods). The bright emission indicates the generation of a similar defect complex due to localized irradiation as in hBN-1. A zoomed-in PL map around the irradiated spot reveals the isolated diffraction-limited bright spot identical to what is observed in hBN-1, which presents stable quantum emission. However, in the subsequent PL map, it is evident that the emitter eventually bleached out, as shown in Figure 3(g). Furthermore, investigation reveals PL count drops to the detector noise level within 10 s under pulse excitation, as shown in Figure 3(h). This instability is observed in almost every studied emitter in hBN-2 crystal (see Supplementary Section S5). This could be attributed to the unstable charge carriers in the excited state, potentially due to the material’s intrinsic doping, which depends on the growth conditions.
@@
In addition to the hBN, our irradiation experiment on SiC (4H-semi insulating) and GaN do not activate any emitters or emitters ensembles as evident by PL map (see Supplementary Section S6). We found that this is independent of the electron beam dose and electron accelerating voltage. We can draw already two conclusions from our experiments: (i) the electron irradiation process works very well to generate single photon emitters in hBN but not in other wide bandgap materials. This indicates that the hBN defects are likely lattice defects and not deposited complexes on the surface, at least using this fabrication method. If latter was the case, this should also work on other materials as a substrate. And (ii), while we could create single photon emitters in both hBN samples, there seems to be a different intrinsic defect density that instabilizes the defects in hBN-2.
IV.2 Theoretical calculations